# SPHERE: Spherical partitioning for large-scale routing optimization

### v1

@@ 4.3 Recursive spherical partitioning
The following algorithm extends the partition cut from Algorithm 1 into a recursive partitioning routine. Its goal is to partition a global $s\!\to\!t$ routing task into a set of smaller subproblems, each confined to an induced sphere whose radius does not exceed a user-defined bound $R_{\max}$. These subproblems are returned as task triples $(H,u,w)$, meaning: “route from $u$ to $w$ within the induced subgraph HH.”
@@
2     radius cap $R_{\max}$, rules $\mathrm{staru},\mathrm{decru},\mathrm{anru}$
@@ 13  $\mathcal{T}\leftarrow\mathcal{T}\cup\{(G^{\mathrm{ind}}_{R_{t}}(t),\,a,\,t)\}$
Figure 3: Recursive splitting. Any side above the radius cap is split again by Algorithm 1; leaves are independent subproblems.
Figure 3 illustrates this recursive structure. Each partition cut splits the problem into two boxes, one for the start-side and one for the target-side. If the radius on a side is within the threshold, the box is recorded as a leaf; otherwise, it is expanded by another partition cut.
@@ 4.3.1 Baseline application
The recursive partition in Algorithm 2 produces a set of leaf tasks, each represented by a triple $(H,u,w)$ where $H$ is an induced spherical subgraph and $u,w$ are its boundary terminals. At this stage, the global $s\to t$ routing problem has been reduced to a collection of independent local subproblems, each small enough to be solved efficiently by standard methods.
@@
2:     radius cap $R_{\max}$, solver $\mathrm{optimize}(\cdot)$
@@
Algorithm 3 describes how to turn the recursive spherical partitioning into a full routing solution. First, the input graph $G$ with terminals $s,t$ is partitioned into a collection of smaller subproblems (line 3). Each subproblem $(H,u,w)$ is an induced subgraph $H$ together with its local terminals $u$ and $w$. Then, for each subproblem, an arbitrary downstream solver $\mathrm{optimize}(H,u,w)$ is applied to compute a local path (lines 4-5). These local solutions are concatenated in the order determined by the anchors of the partition, resulting in a valid global route $\mathcal{R}$ from $s$ to $t$ (line 6). Since the subproblems are independent, they can also be solved in parallel, and the procedure remains agnostic to the choice of solver.
@@ 5.1 Dataset
For our empirical evaluation we use the well-known West-USA-road benchmark suite from the 9th DIMACS Implementation Challenge on Shortest Paths (DIM, 2006). These graphs encode real-world road networks at the level of individual intersections and road segments. Each node corresponds to a road intersection or endpoint, and each edge represents a street segment with a positive weight equal to its physical length. The dataset is widely used in the routing literature as a benchmark for large-scale shortest-path computations, since it combines realistic geometry, heterogeneous edge weights, and large graph sizes (millions of nodes and edges).
@@ 5.2 Experimental setup
For each problem seed $p\in\{1,\dots,30\}$ we draw a source-target pair $(s,t)$ uniformly at random from the node set (rejecting $s=t$). For each $(s,t)$, we run five independent trials with inner seeds $q\in\{1,\dots,5\}$ to capture variability from randomized components. For instance, the anchors are selected randomly according to Definition 6. We report results per instance and aggregate by averaging across the $5$ inner seeds for each problem seed.
Given $(s,t)$, we apply Algorithm 2 and selects an anchor node uniformly at random from this set of nodes (the randomness is controlled by the inner seed $q$). The graph is partitioned into subgraphs around $s$ and $t$ until a maximum radius threshold is met. $R_{max}$ from Algortihm 3 is set to $1800$. Throughout development we found a robust rule of thumb: when routing within a spherical subgraph of hop radius $R$, choosing $R_{\max}\in[\,R/2,\;R/1.7\,]$ delivers high-quality solutions at both low computational cost and effort. For the West-USA instances, a single sphere with $R\approx 3200$ covers the region; thus $R_{\max}=1800$ lies squarely in this band, balancing overlap likelihood and leaf size. Each subgraph shortest-path subproblem is solved with Dijkstra’s algorithm. Because our spherical partitioning produces independent induced subgraphs, we parallelize these subgraph solves and report end-to-end wall-clock time (including partition cut and final concatenation). By contrast, the baselines do not naturally partition into independent leaf tasks so we do not apply the same solving subproblem parallelization to them.
Baselines.
As baselines, we first include an optimal reference obtained by running Dijkstra’s algorithm directly on the full graph. To compare against partition-based approaches, we also evaluate a corridor-routing strategy in which the graph is partitioned into $k=64$ cells using PyMetis; a path is computed on the resulting quotient graph, and Dijkstra’s algorithm is then applied within the induced corridor to recover a detailed route (Karypis and Kumar, 1998). In addition, we test a partition heuristic based on community detection via the Louvain method: routing is first performed on the community graph, and the solution is subsequently refined by applying Dijkstra on the induced subgraph spanned by the relevant communities (default parameters; seeded where applicable) (Blondel et al., 2008). The corridor-routing and Louvain pipelines introduce partitions, yet can not be solved independently and not conditioned on the specific $(s,t)$, so they likewise do not yield independent leaves that one could solve with a plug-in solver and then concatenate repair-free (e.g. fixing feasibility) (Karypis and Kumar, 1998; Blondel et al., 2008).
Metrics. Let $m\in\{\textsc{Sphere},\ \textsc{Corridor},\ \textsc{Louvain}\}$ denote the method. For each problem seed $p$ and inner seed $q$, let $C_{p,q}^{(m)}$ be the route cost returned by $m$, and let $C_{p}^{\star}$ be the shortest-path cost from Dijkstra’s algorithm on the full graph for the same $(s,t)$. Define the relative optimality gap
@@ 5.3 Results
Across 30 West-USA instances with five inner seeds each, SPHERE achieves faster end-to-end runtimes than the three baselines and matches near-optimal route quality. We use performance and accuracy profiles for visualization (Dolan and Moré, 2002; Beiranvand et al., 2017). Both are computed per instance from the median over seeds. Mean-based profiles for runtime and accuracy are presented as well.
@@
Route quality.
Figures 4(b) and 4(d) reports the fraction of instances whose relative optimality gap is at most $\tau$. Dijkstra attains zero gap by construction. SPHERE’s curve dominates the partition baselines across $\tau$, yielding smaller or equal gaps on the majority of instances. This confirms that cutting inside the $s$-$t$ overlap and solving on induced spheres preserves high-quality routes without boundary repair.
Pareto dominance
@@
Robustness and aggregation
Repeating the analysis with means over seeds yields profiles that do not differ materially from the median-based plots, supporting robustness to the randomized anchor selection. Additionally, the full results in the appendix show low dispersion across seeds. For SPHERE, the optimality-gap standard deviation is 0 on 16 instances and is the lowest among all methods on all but one instance. For further Information, we refer to Appendix A.3.
@@ 6 Conclusion and Outlook
We introduced SPHERE, a routing scheme that splits an sstt query by cutting inside the last nonempty overlap of hop-spheres around $s$ and $t$. This placement preserves feasibility without boundary repair, confines each side to an induced subgraph, and allows the two subproblems to be solved independently with standard shortest-path routines and in parallel.
On West-USA-road graphs, SPHERE achieves faster end-to-end times and equal or smaller optimality gaps than a Louvain-based route and a METIS-corridor pipeline on the large majority of instances; runtimes are also more predictable, as heavy-tail slowdowns seen for the baselines largely disappear. In contrast to bidirectional search, which still explores a single coupled state space, SPHERE yields separable leaves. In contrast to static partitions whose cuts ignore $(s,t)$ and can sever bottlenecks, cutting within the overlap keeps the route intact and avoids corridor repair.
Looking ahead, the main levers are the starting rule, the decrement rule, and the anchor selection. Even simple choices already perform well, but data-driven or adaptive variants should improve robustness across graphs and query distributions. A second priority is to tune the radius cap $R_{\max}$ automatically (per instance or region) to balance overlap likelihood against subproblem size. These steps retain the current simplicity and parallelism while targeting further speed and stability.
@@ A.2 Parameters
A.3 Performance per instance

### v2

@@ 4.3 Recursive spherical partitioning
The following procedure extends the partition cut of Algorithm 1 into a recursive framework. The objective is to decompose a global $s\to t$ routing task into a collection of manageable subproblems, each strictly confined to an induced spherical subgraph where the radius does not exceed a bound $R_{\max}$. For a comprehensive overview of all configurable parameters, please refer to Appendix A.2.
Each subproblem is formalized as a task triple $(H,u,w)$, representing a localized search from node $u$ to node $w$ within the induced subgraph $H$. We treat the resulting task set $\mathcal{T}$ as a sequence ordered along the $s\to t$ trajectory. This ordering ensures that concatenating the locally optimal paths induces a valid global path. For simplicity, we refer to this locally optimal path as the path returned by Algorithm 2.
@@
2     bound $R_{\max}$, rules $\mathrm{staru},\mathrm{decru},\mathrm{anru}$
@@ 13  $\mathcal{T}\leftarrow\mathcal{T}\cup\{(G^{\mathrm{ind}}_{R_{t}}(t),\,a,\,t)\}$
Figure 3: Recursive splitting. Any side above the bound is split again by Algorithm 1; leaves are independent subproblems.
Figure 3 illustrates this recursive structure. Each partition cut splits the problem into two boxes, one for the start-side and one for the target-side. If the radius on a side is within the bound, the box is recorded as a leaf; otherwise, it is expanded by another partition cut.
@@ 4.3.1 Baseline application
The recursive decomposition defined in Algorithm 2 yields a set of leaf tasks, each formalized as $(H,u,w)$. This process effectively reduces the high-complexity global $s\to t$ routing problem into a collection of independent, localized subproblems. Because these subproblems are spatially confined and bounded by the bound $R_{\max}$, they are sufficiently small to be solved efficiently using standard routing methods or off-the-shelf optimizers.
@@
2:     bound $R_{\max}$, solver $\mathrm{optimize}(\cdot)$
@@
Algorithm 3 details the integration of recursive spherical partitioning into a complete routing pipeline. The procedure begins by decomposing the input graph $G$ and terminals $(s,t)$ into a sequence of smaller, manageable subproblems with resulting task $(H,u,w)$.
Subsequently, an arbitrary downstream solver, $\mathrm{optimize}(H,u,w)$, is applied to each subproblem to compute a local path. These local solutions are then concatenated in the order established by the partition’s anchors, forming a globally valid $s\to t$ route $\mathcal{R}$. A critical advantage of this framework is that because the subproblems are mutually independent, they can be solved in parallel. Furthermore, the procedure remains strictly solver-agnostic, allowing the utilization of various downstream optimizer without modification to the underlying partitioning logic.
4.4 Optimality Conditions
[…]
In the following, we demonstrate that Algorithm 1, when applied to unweighted graphs using the monotone decrement rule from Example 4.6, preserves optimality. We thus provide a stronger result compared to the the general weighted setting, where Proposition 4.9 ensures the preservation of route feasibility only.
Furthermore, this result allows to compute a globally shortest path without invoking any additional optimization routines or solver-specific procedures, provided we apply the recursion until the radii are equal to $1$.
We begin by formalizing the Strong Minimality Condition, which characterizes the minimal non-empty overlap induced by the monotone decrement rule in terms of the associated radius tuple $(R_{s},R_{t})$.
Definition 4.11 (Strong Minimality Condition).
A tuple $(R_{s},R_{t})$ is strongly minimal if
$O(R_{s},R_{t})\neq\emptyset,$
$O(R_{s}-1,R_{t})=\emptyset,$
$O(R_{s},R_{t}-1)=\emptyset.$
The next result shows by combining Definitions 4.4 and 4.11 that strong minimality forces all overlap vertices to lie simultaneously on the boundaries of both spheres determined by $(R_{s},R_{t})$.
Lemma 4.12 (Boundary property).
If $(R_{s},R_{t})$ is strongly minimal, then for every $v\in O(R_{s},R_{t})$,
$d(s,v)=R_{s}\quad\text{and}\quad d(t,v)=R_{t}.$
Proof 4.13.
Let $v\in O(R_{s},R_{t})$. If $d(s,v)\leq R_{s}-1$, then $v\in O(R_{s}-1,R_{t})$, contradicting strong minimality. Hence $d(s,v)=R_{s}$. The argumentation for $t$ is analogous, yielding $d(t,v)=R_{t}$.
Having established the boundary property of strongly minimal overlaps, we now derive the key consequence of the monotone decrement rule. The following Meet-in-the-middle equality formalizes that a strongly minimal radius tuple satisfies
$R_{s}+R_{t}=d(s,t),$
thereby identifying the last overlap of two spheres.
Lemma 4.14 (Meet-in-the-middle equality).
If $(R_{s},R_{t})$ is strongly minimal, then
$R_{s}+R_{t}=d(s,t).$
Proof 4.15.
Choose $a\in O(R_{s},R_{t})$. Since $a\in S_{R_{s}}(s)$ and $a\in S_{R_{t}}(t)$, there exists an sstt path via $a$ of length at most $R_{s}+R_{t}$, hence $d(s,t)\leq R_{s}+R_{t}$.
Assume for contradiction that $d(s,t)\leq R_{s}+R_{t}-1$. Let $P$ be a shortest sstt path of length $L=d(s,t)$. Let $v$ be the vertex on $P$ at distance $R_{s}$ from $s$. Since $P$ is the shortest sstt path, then $d(s,v)=R_{s}$ and $d(t,v)=L-R_{s}\leq R_{t}-1$. Thus $v\in O(R_{s},R_{t}-1)$ contradicts strong minimality. Therefore $d(s,t)\geq R_{s}+R_{t}$, and equality follows.
Lemma 4.14 constitutes the central argument of our main optimality theorem: it ensures that any anchor selected from the minimal overlap lies on a globally shortest path and thus enables recursive preservation of optimality. Note that all definitions and lemmas apply to any terminal pair $(u,w)$ (replacing $(s,t)$), in every recursive subproblem.
Theorem 4.16 (Global preservation).
Let $G=(V,E)$ be a finite, connected, undirected, and unweighted graph and assume that the monotone decrement rule is used. Then the concatenated sstt path returned by Algorithm 2 is a globally shortest path in $G$.
Proof 4.17.
We argue by induction over the recursion tree:
Local optimal split. Fix any subproblem with endpoints $(u,w)$ and a strongly minimal pair $(R_{u},R_{w})$ as induced by undirected graphs and the monotone decrement rule with anchor $a\in O(R_{u},R_{w})$. Applying Lemma 4.12, we observe that the anchor $a$ lies on the boundaries of both spheres, such that $d(u,a)=R_{u}$ and $d(a,w)=R_{w}$. It follows from Lemma 4.14 that these radii sum exactly to the distance between $u$ and $w$, i.e. $R_{u}+R_{w}=d(u,w)$. Hence
$d(u,a)+d(a,w)=d(u,w),$
so $a$ lies on a shortest uuww path and the split $(u,w)\mapsto(u,a),(a,w)$ is consistent with global optimality for that subproblem.
Inductive step. Assume recursively computed solutions are shortest for the two child subproblems $(u,a)$ and $(a,w)$. Their concatenation has length $d(u,a)+d(a,w)=d(u,w)$, hence is shortest for the parent subproblem. Applying this from leaves to the root yields a globally shortest sstt path.
@@ 5.1 Dataset
For our empirical evaluation, we use the West-USA-road benchmark suite from the 9th DIMACS Implementation Challenge on Shortest Paths [DIMACS2006SP]. This dataset represents real-world road networks. Nodes correspond to intersections and edges represent road segments with positive weights equal to physical length.
@@ 5.2 Experimental Setup
For each problem instance $p$ (with according instance seed), we draw a source-target pair $(s,t)$ uniformly at random. To capture variability from the randomized anchor selection, we run five independent trials with inner seeds $q$ for each source-target pair. We report results aggregated by the median and mean across these seeds. Further details regarding graph dimensions and seed selection are provided in Appendix A.3.
We apply Algorithm 2 with a maximum radius bound of $R_{\max}=1800$ (see Appendix A.3 for tuning rationale). Because SPHERE generates independent induced subgraphs, we solve the resulting leaf tasks in parallel using Dijkstra’s algorithm and report the end-to-end wall-clock time, including the partitioning and concatenation overhead. [^1] The graph data and implementation will be made publicly available upon the manuscript’s acceptance.
We compare SPHERE against an optimal reference using Dijkstra’s algorithm on the full graph, as well as Louvain and Corridor routing. Unlike SPHERE, these static partitioning schemes compute cuts globally rather than conditioning on the specific $(s,t)$ query. Consequently, they cannot guarantee feasibility without boundary repair, nor do they naturally yield the mathematically independent subproblems required for the same drop-in parallel execution framework applied to our method. While static partitions could theoretically be parallelized, they often require shared boundary synchronization or global quotient graphs to ensure path consistency, whereas SPHERE subproblems are self-contained by construction. (See Appendix A.3 for implementation and library details).
Let $m\in\{\emph{SPHERE},\ \textsc{Corridor},\ \textsc{Louvain}\}$ denote the method. For each problem seed $p$ and inner seed $q$, let $C_{p,q}^{(m)}$ be the route cost returned by $m$, and let $C_{p}^{\star}$ be the shortest-path cost from Dijkstra’s algorithm on the full graph for the same $(s,t)$. Define the relative optimality gap
@@ 5.3 Results
Across 30 West-USA instances with five inner seeds each, SPHERE achieves faster end-to-end runtimes than the three baselines and matches near-optimal route quality. We use performance and accuracy profiles for visualization [dolan2002benchmarking, beiranvand2017best]. Both are computed per instance from the median over seeds. Mean-based profiles for runtime and accuracy are presented as well.
@@
Figures 4(b) and 4(d) report the fraction of instances whose relative optimality gap is at most $\tau$. Dijkstra attains zero gap by construction. SPHERE’s curve dominates the partition baselines across $\tau$, yielding smaller or equal gaps on the majority of instances. This confirms that cutting inside the sstt overlap and solving on induced spheres preserves high-quality routes without boundary repair.
@@
Repeating the analysis with means over seeds yields profiles that do not differ materially from the median-based plots, supporting robustness to the randomized anchor selection. Additionally, the full results in Appendix A.4 show low dispersion across seeds. The results indicate reliability across varying query geometries and suggest that spherical decomposition effectively captures the core topological features necessary for high-quality routing, regardless of internal stochasticity. For SPHERE, the optimality-gap standard deviation is 0 on 16 instances and is the lowest among all methods on all but one instance. For further information, we refer to Appendix A.4.
@@ 6 Conclusion and Outlook
We introduced SPHERE, a routing scheme that splits an sstt query by cutting inside the last nonempty overlap of hop-spheres around $s$ and $t$. This placement preserves feasibility without boundary repair, confines each side to an induced subgraph, and allows the subproblems to be solved independently with standard shortest-path routines and in parallel. On million-scale road networks, SPHERE achieves faster runtimes and smaller optimality gaps than state-of-the-art static and community-based partitioning methods. Notably, our approach eliminates the heavy-tail runtime outliers common in exact solvers, providing high statistical stability across randomized anchor selections.
Future work will explore data-driven anchor selection to further improve robustness across diverse graph topologies. Additionally, SPHERE can serve as a wrapper for complex neural combinatorial optimizers that typically struggle with large-scale instances. Another direction could be to automate the tuning of the radius cap $R_{\max}$ to adaptively balance computational budgets against query complexity. These advancements will maintain SPHERE’s simplicity and parallelism while further enhancing speed and stability.
@@ A.2 Parameters
A.3 Implementation Details and Hyperparameters
Graph Dimensions.
The West-USA-road network consists of $6,262,104$ nodes and $15,248,146$ edges. The large scale of this network makes it a challenging benchmark for standard exact solvers.
Instance and Seed Selection.
The $30$ problem instances $p$ are defined by fixed global seeds to ensure reproducibility of the $(s,t)$ pairs. For each $p$, 55 inner seeds $q$ control the stochasticity of the $\mathrm{anru}$ function in Algorithm 1.
Radius bound ($R_{\max}$).
The tunable bound $R_{\max}$ controls the maximum diameter of the subproblems generated by Algorithm 3. Throughout development, we established a robust rule of thumb: when routing within a spherical subgraph of hop radius $R$, choosing $R_{\max}\in[\,R/2,\;R/1.7\,]$ delivers high-quality solutions at both low computational cost and effort. For the West-USA benchmark instances, a single sphere with $R\approx 3200$ generally covers the necessary region. Thus, our chosen bound of $R_{\max}=1800$ lies squarely in this optimal band, carefully balancing the likelihood of an sstt overlap against the complexity of the resulting leaf tasks.
Baseline Configurations.
The baseline pipelines were implemented using standard libraries to ensure a fair comparison. The corridor-routing baseline utilizes PyMetis to compute the initial $k=64$ cell partition. The Louvain community-detection baseline was run using its default modularity-optimization parameters, seeded where applicable to ensure deterministic execution. It is worth noting that because these static partitioning schemes compute cuts globally rather than conditioning on the specific $(s,t)$ query, they cannot guarantee feasibility without boundary repair, nor do they yield strictly independent leaves suitable for the same parallel execution framework applied to SPHERE.
A.4 Performance per instance
Continued on next page.